OpenStudy (anonymous):

decompose each expression into partial fraction -x+5/(x-1)(x+1)

OpenStudy (anonymous):

@jdoe0001

OpenStudy (turingtest):

how far have you gotten?

OpenStudy (turingtest):

btw is this (-x+5)/(x-1)(x+1) ?

OpenStudy (anonymous):

i have no idea how to this probelm can you show me step by step

OpenStudy (turingtest):

did I type the problem correctly?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

reference this http://tutorial.math.lamar.edu/Classes/Alg/PartialFractions.aspx I'm a bit busy, so i'll give you fairly minimal info, sorry

OpenStudy (turingtest):

start of by breakingit up into \[{x-5\over(x-1)(x+1)}=\frac A{x-1}+\frac B{x+1}\]then multiply both sides by \[(x-1)(x+1)\]and tell me what you get